40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 25 23:15:11 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking fpmibd.pdf
    Reply  Wed Sep 25 23:59:29 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking XARM_OLTF.pdfYARM_OLTF.pdfXARM_ITMXresponse.pngYARM_ITMYresponse.pngfree_running.pdf
       Reply  Thu Sep 26 01:49:28 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking MICH_OLTF.pdfBS-RS55Q.pngETMX-RS55.pngETMY-RS55.pngplot.pdf
       Reply  Thu Sep 26 23:02:40 2013, rana, Update, LSC, FPMI noise caused by ARM locking Xarm_sweep_130926.pdflsc.pdferr.png
          Reply  Fri Sep 27 00:48:53 2013, Masayuki, Update, LSC, FPMI noise caused by ARM locking phase_badget.pngFPMI_noises.png
Message ID: 9168     Entry time: Fri Sep 27 00:48:53 2013     In reply to: 9167
Author: Masayuki 
Type: Update 
Category: LSC 
Subject: FPMI noise caused by ARM locking 


Hidden in Nakano-kun's previous entries was that the phase margin of the X-Arm was only 9 degrees!! This extremely close to instability and makes for huge gain peaking. The feedback loop is increasing noise above 100 Hz rather than suppress. After some tweaks of the LSC filters we got a much more stable loop/.

So we today started to examine the sources of phase lag in the arm cavity sweeps. There were a few unfortunate choices in the XARM LSC filter bank which we tuned to get less delay.

Then I wrote a bunch of detail about how that worked, but the ELOG ate my entry because it couldn't handle converting my error signal noise plot into a thumbnail. Then it crashed and I restarted it. We also have now propagated the changes to the Y arm by copy/paste the filters and the result there is pretty much the same: low phase margin is now 38 deg phase margin. Noise is less bad.


[Rana, Masayuki

 I made the plot of the phase of the digital filters which Rana change and also  of the AA, AI, DAA, DAI filters. Now the biggest phase delay come from the timedelay of the digital system.


The UGF is around 150 Hz at that frequency the time delay has biggest phase  delay. Second one is the FM9 filter (this filter is BOOST filter). Then we have the AA filter, AI filter and so on, but these delay is roughly 5 degree.

As I said in previous entry, the time delay of the XARM control is roughly 300 usec, and we have 120 usec even only in C1SUS. Also between the C!SUS and C1LSC we have another 120 usec time delay. We want to increase the UGF to 300 Hz but because of the time delay of the digital system we cannot increase. So we should fix this problem.


After changing these filters, the FPMI noise is become better at high frequency. Before we have peak around the 100 Hz (because of 8 degree phase margin...), but they are gone. i attached the noise spectrum. This plot is measured by the real time calibration output. But even then, you can see the extra noise around 100 Hz in FPMI conpare to only MICH.




ELOG V3.1.3-